5. Independent events A and B are such that P(A) = 0.45 and P (A B) = 0.18. Find:

a) P(B)

b) P(A U B)

c) P (A, λ B)
Answer #1

Q.5) Given that, A and B are independent event such that,

P(A) = 0.45 and P(AnB) 0.18

a) since, A and B are independent event,

ΡΑλ B) = P(A) * P(B)

=> 0.18 = 0.45 * P(B)

> P(B) = 0.18 0.45

> P(B) = 0.40

b)

P(AU B) = P(A) + P(B) _ P(An B)

> P(A U B) = 0.45 0.40-0.18

=> P(A U B) = 0.67

c)

P(A' cap B') = P(A cup B)'

=> P(A' cap B') = 1-P(A cup B)

=> P(A' cap B') = 1- 0.67

=> P(A, λ B) 0.33